OpenStudy (anonymous):

Electrical Current. The current I in an electrical conductor varies inversely as the resistance R of the conductor. The current is 1/7 ampere when the resistance is 14406 ohms. What is the current when the resistance is 1176 ohms?

OpenStudy (anonymous):

V=I R

OpenStudy (anonymous):

I=V/R

OpenStudy (anonymous):

1/7 = V/14406 V=2058 use that to find second part I= 2058/1176